A pipe bending machine used to produce ship system components was operated as follows:

  • The period of operation was 25 days @ 10 hrs per day.
  • It produced 800 total pipe details, of which 40 were rejected by quality inspectors.
  • The total setup time for the machine during the 25 days was 22 hours.
  • The machine had 3 unplanned breakdown events
    • 1.5 hours for a power failure
    • 0.5 hours for computer network failure
    • 2.2 hours for lubricant pump failure
  • The theoretical cycle time for the parts averages 10 minutes per part.
  • 30 minutes of planned downtime was experienced when workers attended a Labor Union Meeting

(show calculations)

         a. What is the overall equipment efficiency (OEE) value? (show work)

         b. What are the quality; performance, and availability rates? (show work)

         c. Between Availability, Quality and Performance, which measure would you

             prioritize when beginning improvement efforts?
Answer #1

OEE - Overall Equipment Effectiveness relies on 3 factors - Availability, Performance, and Quality.

i.e. OEE=Availability*Performance*Quality

Run Time = Planned Production Time − Stop Time = ((25*10)-0.5-22) - (1.5+0.5+22) = 203.5 hrs.

Availability is the takes into account the unplanned and planned stoppages

Availability =  Run Time / Planned Production Time
= 203.5/ (25*10)
= 0.814

i.e. Availability is 81.4%

Performance takes into account the slow cycle time.

Performance = (Ideal Cycle Time × Total Count) / Run Time
= (1/6 * 800)/ 203.5 Cycle time is 10mins per part i.e. 1/6 hr.
= 0.655

i.e. Performance is 65.5%  

Quality is regarding the defects(or errors) created or done  in the process

Quality = Good Count / Total Count = (800-40)/800 since, 40 out of 800 details are defective
= 0.95

i.e. Quality is 95%

Hence, OEE = 81.4%*65.5%*95% *100
= 50.65%

For improvement in the above mentioned case, priority to be given to improving the Performance of the equipment as there is a huge scope for improving this parameter.